OpenStudy (anonymous):

Find the distance between the points (3, 0) and (-1, 5).

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

OpenStudy (anonymous):

3-1)2+0-5)2

hartnn (hartnn):

3-(-1) = 3+1 = 4 4^2 = 16 ok?

OpenStudy (anonymous):

29

hartnn (hartnn):

how ?

hartnn (hartnn):

5 ^2 = 25

OpenStudy (anonymous):

right and i so 4+25=

hartnn (hartnn):

where did 4 come from ????

OpenStudy (anonymous):

3-1=2

hartnn (hartnn):

3-(-1) = 4 4^2 =16

OpenStudy (anonymous):

16+25

OpenStudy (anonymous):

so 41

hartnn (hartnn):

yes, 41, now taking square root is remaining.

hartnn (hartnn):

keep it as \(\sqrt41\) or calculate it using calculator

hartnn (hartnn):

u there ?

OpenStudy (anonymous):

Formula is right,t, wrong steps though

hero (hero):

Why can't one just simply do this: